OpenStudy (anonymous):

simplify the quotient? sqrt3x / sqrt18

OpenStudy (anonymous):

answer or explanation

OpenStudy (anonymous):

both ?

OpenStudy (anonymous):

\[\frac{ \sqrt{3x} }{ \sqrt{18} } = \frac{ \sqrt{54x} }{ 18 } = \frac{ 3\sqrt{6x} }{ 18 } = \frac{ \sqrt{6x} }{ 6 }\]

OpenStudy (anonymous):

so first multiply by sqrt(18) / sqrt(18) to get rid of the sqrt on the bottom... Then simplify sqrt(54x)...Then cancel out the 3 and replace the 18 on the bottom with 6

OpenStudy (anonymous):

hmm you could also multiply top and bottom by \(\sqrt2\) you get \[\frac{\sqrt{3x}}{\sqrt{18}}\times \frac{\sqrt2}{\sqrt2}=\frac{\sqrt{6x}}{\sqrt{36}}=\frac{\sqrt{6x}}{6}\]

OpenStudy (anonymous):

thanks somuch! can you help with another?